OpenStudy (anonymous):

URGENT The range of the function f(x) = x + 5 is {7, 9}. What is the function’s domain? A) {2, 4} B) {-2, -4} C) {12, 14} D) {-12, -14} E) {0, 5}

OpenStudy (anonymous):

The range is the f(x) or y value so you need to solve these equations to find the domain. \[x+5=7\] \[x+5=9\]

OpenStudy (anonymous):

so is {2,4}????? @peachpi

OpenStudy (texaschic101):

first, sub in 7 for x in f(x) = x + 5

OpenStudy (texaschic101):

f(x) represents y...and the y values are your domain. so what is y in this first equation ? f(x) = 7 + 5......is the same as y = 7 + 5... do the same for the other one... f(x) = x + 5....when x = 9...and remember f(x) = y.. y = 9 + 5...so what is y ?

OpenStudy (anonymous):

now I understand!! so the anwser is {12,14}??? @texaschic101

OpenStudy (anonymous):

it's {2,4}